OpenStudy (anonymous):

What is the joint probability P(x>y)of F(x,y)=xy where 0<=x<=2 and 0<=y<=1??? I keep getting 1/8 while the answer key says 7/8

OpenStudy (anonymous):

correct me if i am wrong, but do you compute via \[\int_0^1\int _0^x xydxdy\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

dydx.. I think

OpenStudy (anonymous):

yeah that would make more sense

OpenStudy (anonymous):

hmm in that case you do get \(\frac{1}{8}\) don't you? i wonder why it is wrong

OpenStudy (anonymous):

it is \(P(X>Y)\) right? not the other way around

OpenStudy (anonymous):

Okay, imma go with my gut. I guess the answer key is wrong. Thank You

OpenStudy (anonymous):

hello joe!!!

OpenStudy (anonymous):

Your sample space looks like this: So your integration limits should be:\[\int\limits_{0}^1\int\limits_{y}^2xydxdy\]i think.